OpenStudy (anonymous):

A certain kind of candy comes in mixed bags of five shapes: 40% are flowers, 30% are snowflakes, 20% are butterflies, and 10% are diamonds. Ethan's favorites are the snowflakes and the diamonds. What is the probability that the next one he picks (at random) will be one of his favorites?

OpenStudy (zarkon):

add snowflakes and diamonds

OpenStudy (anonymous):

I did I got 40/100 but it's saying it's incorrect. I've tried reducing it to 1/25 and 1/50 but it still says it's incorrect.

OpenStudy (zarkon):

40/100=4/10=2/5

OpenStudy (anonymous):

Oh, thanks! How did you get 4/10?

OpenStudy (zarkon):

\[\frac{40}{100}=\frac{4\times 10}{10\times 10}=\frac{4\times \cancel{10}}{10\times \cancel{10}}=\frac{4}{10}\]

OpenStudy (anonymous):

Ahh, okay! I get it. Thank you so much!! :)